The oxidation of 25 mL of a solution containing Fe2+ requires 29 mL of 0.0230 M K2Cr2O7 in acidic solution. Balance the following equation. Do NOT include the states of each species. Cr2O72− + Fe2+ + H+ → Cr3+ + Fe3+ → Calculate the molar concentration of Fe2+: M
Answer #1

Answer

0.1601M

Explanation

The balanced chemical equation is

Cr2O72- + 6Fe2+ + 14H+ --------> 2Cr3+ + 6Fe3+ + 7H2O

Stoichiometrically , 1mole of Cr2O72- react with 6moles of Fe2+

Moles of Cr2O72- consumed = (0.0230mol/1000ml) × 29.0ml = 0.000667mol

moles of Fe2+ present in the solution = 6 × 0.000667mol =0.004002mol

Molar concentration of Fe2+ = ( 0.004002mol/25ml ) ×1000ml = 0.1601M